a coin collector has $2.25 in nickels and dimes. She has 3 more dimes than nickels how many nickels does she have?
Answer by mananth(16946)

a coin collector has $2.25 in nickels and dimes. She has 3 more dimes than nickels how many nickels does she have?
..
let nickels be x
dimes = x+3
..
5x+10(x+3)=225
5x+10x+30=225
15x+30=225
add -30
15x=225-30
15x=195
x=195/15
x=13 nickels
dimes = 16.
...
16*10+13*5=225
...
m.ananth@hotmail.ca